# PRISM-DP: Spatial Pose-based Observations for Diffusion-Policies via Segmentation, Mesh Generation, and Pose Tracking

### v2

@@ Abstract
Diffusion-based visuomotor policies generate robot motions by learning to denoise action-space trajectories conditioned on observations. These observations are commonly streams of RGB images, whose high dimensionality includes substantial task-irrelevant information, requiring large models to extract relevant patterns. In contrast, using more structured observations, such as the spatial poses (positions and orientations) of key objects over time, enables training more compact policies that can recognize relevant patterns with fewer parameters. However, obtaining accurate object poses in open-set, real-world environments remains challenging. For instance, it is impractical to assume that all relevant objects are equipped with markers, and recent learning-based 6D pose estimation and tracking methods often depend on pre-scanned object meshes, requiring manual reconstruction. In this work, we propose PRISM-DP, an approach that leverages segmentation, mesh generation, pose estimation, and pose tracking models to enable compact diffusion policy learning directly from the spatial poses of task-relevant objects. Crucially, because PRISM-DP uses a mesh generation model, it eliminates the need for manual mesh processing or creation, improving scalability and usability in open-set, real-world environments. Experiments across a range of tasks in both simulation and real-world settings show that PRISM-DP outperforms high-dimensional image-based diffusion policies and achieves performance comparable to policies trained with ground-truth state information. We conclude with a discussion of the broader implications and limitations of our approach.
Keywords: Imitation Learning, 6D Pose Estimation and Tracking, Manipulation
1 Introduction
@@
To address these challenges, we introduce Pose-tRacking via Image Segmentation and Mesh-generation Diffusion Policies (PRISM-DP), an approach that leverages segmentation, mesh generation, pose estimation, and pose tracking models to enable compact diffusion policy learning directly from the spatial poses of task-relevant objects. More specifically, PRISM-DP applies a segmentation model to isolate task-relevant objects in input images, it uses the resulting segmentation masks to automatically create 3D triangulated meshes for all target objects using a mesh generation model, then it uses the newly created meshes to initialize a pose estimation and tracking model. Subsequently, the diffusion policy is conditioned on both the robot state and the continuously estimated object poses for the target objects. Crucially, because PRISM-DP uses a mesh generation model, it eliminates the need for manual mesh processing or creation, improving scalability and usability in open-set, real-world environments.
@@
Through several experiments in both simulated and real-world settings, we demonstrate that PRISM-DP facilitates the learning of compact diffusion policies that achieve performance comparable to policies trained using ground-truth object pose information or manually generated meshes, while significantly surpassing image-based diffusion policies with similar parameter counts. To our knowledge, our work is the first to show that diffusion policies conditioned on pose-based observations can consistently outperform those conditioned on raw image observations in a real-world setting, suggesting exciting new directions for future research. Our implementation is publicly available, aiming to foster future research into scalable policy learning through open-set mesh generation and pose estimation.  [^1] https://github.com/Apollo-Lab-Yale/prism
2 Related Works
@@
6D Pose Estimation and Tracking. 6D pose estimation and tracking are foundational technologies in robotics, with decades of extensive study and advancement [4]. Pose estimation refers to the process of determining the 3D position and orientation (collectively known as the 6D pose) of an object from sensor data, such as images or point clouds. Traditionally, pose estimation and tracking have been treated as distinct problems. Pose estimation methods are broadly categorized into model-based approaches, which rely on known CAD or mesh models of the object [5, 6, 7], and model-free approaches, which instead use a few reference images of the target object [8, 9, 10]. Although model-free methods appear more user-friendly, they often require fine-tuning for each novel object [11], suffer from inefficiencies due to internal reconstruction processes [12], or fail under conditions of occlusion and low texture [13]. As a result, model-based methods are generally preferred for achieving robust and reliable pose estimation.
@@
Pose tracking, by contrast, focuses on continuously updating an object’s 6D pose over time by exploiting temporal information from sequential video frames. It is the task of estimating an object’s changing position and orientation across a sequence of frames, typically with an emphasis on efficiency and smoothness. Like pose estimation, tracking methods can be divided into model-free [14, 15] and model-based approaches [16, 17].
@@
Policy Learning with Poses. Pose information provides a compact and structured description of the environment, making it ideal for policy learning. However, deploying pose-based policies in the real world remains challenging due to the difficulty of acquiring accurate 6D poses.
External tracking methods such as AprilTags [18, 19] or Motion Capture [20, 21] are commonly used but require specialized hardware and controlled environments. Others reconstruct scenes into point clouds or voxel maps using depth sensors [22, 23, 24] or LiDAR [25, 26, 27], but such methods are sensitive to environmental changes and require repeated preprocessing if changes occur. Diffusion policies initially trained with ground-truth poses in simulation have been adapted for real-world tasks via visual encoders that process input images [1, 28, 29], but this technique comes at the cost of three to four times increased model size.
Recent work has incorporated FoundationPose [2] into learning-based robotics pipelines [30, 3]. However, Pan et al. [30] use poses only for task planning, while Hsu et al. [3] requires manual mesh scans using consumer devices. In contrast, our approach removes the need for manual reconstruction by integrating segmentation and mesh generation for open-set 6D pose tracking.
3D Generation. The success of large-scale generative models for text and images [31, 32] has recently extended to 3D AI-Generated Content (3D AIGC), with diffusion [33] and Transformer-based architectures [34] driving progress in mesh generation. Recent methods fall into two categories: reconstruction-based approaches [35, 36, 37], which first generate intermediate 3D representations before converting them into meshes, and autoregressive approaches [38, 39, 40], which directly model mesh structures through sequence learning.
Reconstruction-based methods tend to produce higher visual fidelity [35, 36, 37], while autoregressive methods generate meshes with more concise and elegant topology [38, 39, 40] but often at the cost of visual quality [41]. Given the importance of geometric realism for downstream pose estimation and tracking, our approach uses a reconstruction-based mesh generation approach. Specifically, our current prototype implementation (covered below) uses the Meshy model [41, 42].
3 Technical Overview
@@
(2) A segmentation model produces binary masks for each target object at their respective frames of first visibility: $\mathbf{M}_{i}=\mathrm{Segmentation}(\mathbf{I}_{F_{i}}),\ \forall i\in\{1,...,J\}$. Each pixel in a mask $\mathbf{M}_{i}\in\mathbb{R}^{H\times W}$ is $1$ if part of the $i$-th object at time $t$ and 00 otherwise.
@@
(3) Each mask $\mathbf{M}_{i}$ is passed to a mesh generation model to obtain a triangular mesh: $\mathcal{M}_{i}=\mathrm{MeshGen}(\mathbf{I}_{F_{i}}*\mathbf{M}_{i}),\forall i\in\{1,...,J\}$. The operation $\mathbf{I}_{F_{i}}*\mathbf{M}_{i}$ denotes elementwise multiplication, preserving RGB values in $\mathbf{I}_{F_{i}}$ where the mask is one and setting all other pixels to black (as seen in Fig. 1).
@@
(4) An approximate pose is calculated at the frames of first visibility for all target objects: $\hat{\mathbf{T}}_{i}^{F_{i}}=\mathrm{PoseEstimator}(\mathbf{I}_{F_{i}},\mathbf{D}_{F_{i}},\mathcal{M}_{i},\mathbf{K}),\ \forall i\in\{1,...,J\}$.
@@
(5) The pose estimates at the the frames of first visibility are used to bootstrap a pose tracking model per object: $\hat{\mathbf{T}}_{i}^{t}=\mathrm{PoseTracking}(\hat{\mathbf{T}}_{i}^{t-1},\mathbf{I}_{t},\mathbf{D}_{t},\mathcal{M}_{i},\mathbf{K}),\ \forall i\in\{1,...J\},\ \forall t\in\{F_{i},...T\}$. If an object is not visible at the beginning of a demonstration (i.e., $F_{i}\neq 1$ for some $i$), then the object poses at frames preceding $F_{i}$ are considered null, set as $4\times 4$ matrices of all zeros.
@@
The output of Step 5 is a set of pose estimates over all task-relevant objects and time points. The pose estimates at time $t$ are concatenated with their corresponding robot state, $\mathbf{s}_{t}$, to create a sequence of observations, which are finally used to train a diffusion policy in a compact, low-dimensional space. For interested readers, training model details are found in the Appendix (§A.1).
4 Evaluation
4.1 Prototype System
In this subsection, we provide implementation details for our prototype system that instantiates the approach outlined in §3. For the $\mathrm{Segmentation}$ subroutine, the base model in our system is Segment Anything 2 (SAM2) [43]. Task-relevant objects are specified to SAM2 by clicking on the target objects within the sequence of RGB images on their first visible frames, $\mathbf{I}_{F_{i}}$. Such inputs are called “point prompts”.
Importantly, our current system assumes a human-in-the-loop during task execution. Initially, the poses of all task-relevant objects are set to null values (all zeros). When a target object first appears in the field of view, the user pauses execution and provides point prompts to SAM2 through a GUI. FoundationPose then uses the resulting segmentation mask and corresponding mesh object (cached during training) to initialize the object’s pose via pose estimation. Execution then resumes, and FoundationPose continuously updates this object’s pose through tracking. This initialization process is performed separately for each target object.
4.2 Simulation Experiments
Experimental Settings. To evaluate the effectiveness of PRISM-DP, we conduct extensive experiments in simulation across five tasks from Robosuite [44], using a 7-DOF simulated Franka Panda arm. The control is executed in end-effector space, with rotations represented as quaternions. As illustrated in Fig. 4, the tasks are lifting a red cube (Lift), sorting a can into its designated compartment (Can), inserting a square nut onto a square peg (Square), stacking a red cube onto a green cube (Stack), and placing a mug into a drawer for cleanup (Mug).
The simulation evaluation includes a comprehensive set of baselines and architectural variants, as summarized in Table 1. Evaluated methods include Transformer-based (DP-T) and UNet-based (DP-U) diffusion policies trained with ground-truth object poses (GTP), raw RGB images (Img), and estimated 6D poses using ground-truth meshes (GTM). We also include larger image-based variants, DP-T-L Img and DP-U-L Img, that are 8 times and 11 times larger than their standard counterparts (detailed further in the following paragraph). Our proposed method with a Transformer and a UNet is denoted as PRISM-DP-T and PRISM-DP-U.
@@ 4.2 Simulation Experiments
Figure 4: Simulation tasks in Robosuite. Colored 3D bounding boxes indicate objects whose poses are estimated and tracked by FoundationPose.
Table 1: Success rates (SR) and training time per epoch (TE) across five Robosuite tasks. TE is recorded in seconds. Evaluated methods include Transformer-based (DP-T) and UNet-based (DP-U) diffusion policies trained with ground-truth object poses (GTP), raw RGB images (Img), and estimated 6D poses using ground-truth meshes (GTM). We also include larger image-based variants, DP-T-L Img and DP-U-L Img, that are 8 times and 11 times larger than their standard counterparts. Our proposed method with a Transformer and a UNet is denoted as PRISM-DP-T and PRISM-DP-U.
@@
Results. As shown in Table 1, PRISM-DP consistently outperforms image-conditioned diffusion policies with the same network capacity while having better training efficiency, regardless of whether a Transformer or UNet architecture is used. It achieves comparable performance to diffusion policies trained with poses obtained from ground-truth meshes, indicating that learning-based mesh generation is sufficiently accurate to support downstream pose tracking and policy learning for these tasks. While scaling up the parameter count by 8 times and 11 times in the Transformer-based and UNet-based image models (DP-T-L Img and DP-U-L Img) closes the performance gap for simpler tasks such as Lift and Can, this improvement comes at a significant computational cost and fails to generalize to more challenging tasks. For instance, in tasks like Stack, Square, and Mug, PRISM-DP still outperforms these large image-based baselines, underscoring the value of low-dimensional structured observations.
@@
Analysis on Pose Estimation with Generated Meshes. To further understand the efficacy of PRISM-DP, we evaluate the quality of mesh generation as well as the accuracy of pose estimation and tracking. For mesh generation, we compare meshes produced by the generative model against ground-truth meshes from Robosuite. For pose estimation, we compare the predicted poses, estimated using the generated meshes, to the ground-truth poses transformed into the camera frame.
To assess mesh quality, we render both ground-truth and generated meshes from 30 randomly sampled viewpoints using Blender [48], and compute two commonly used perceptual metrics: Peak Signal-to-Noise Ratio (PSNR) [49] and Fréchet Inception Distance (FID) [50]. PSNR measures pixel-level reconstruction fidelity, where higher values indicate closer alignment, with 20–40 dB considered high quality [51, 52, 53]. FID assesses similarity in feature space, where lower scores are better, and values below 10 typically reflect high-quality generation [54, 51, 52].
@@
Figure 5: Top left: Examples of ground-truth and generated meshes, shown in the upper and lower rows, respectively. Top right: Visualization of estimated and ground-truth poses of one object during a rollout. Bottom: Average PSNR, FID, and the average positional and orientation errors in radian across all objects in simulation training datasets.
@@
4.3 Real-World Experiments
Experimental Settings. To further validate the proposed approach, we conduct real-world experiments using a dual-arm robotic system (Fig. 6). The setup comprises two 7-DOF xArm7 manipulators, each mounted on a 1-DOF linear actuator. One arm is equipped with a RealSense D435i RGB–D camera for viewpoint control, while the other is fitted with a parallel gripper for manipulation. This configuration enables the policy to dynamically select viewpoints during task execution.

### v3

@@ Abstract
Diffusion policies generate robot motions by learning to denoise action-space trajectories conditioned on observations. These observations are commonly streams of RGB images, whose high dimensionality includes substantial task-irrelevant information, requiring large models to extract relevant patterns. In contrast, using structured observations like the spatial poses of key objects enables training more compact policies with fewer parameters. However, obtaining accurate object poses in open-set, real-world environments remains challenging, as 6D pose estimation and tracking methods often depend on markers placed on objects beforehand or pre-scanned object meshes that require manual reconstruction. We propose PRISM-DP, an approach that leverages segmentation, mesh generation, and pose tracking models to enable compact diffusion policy learning directly from the spatial poses of task-relevant objects. Crucially, by using a mesh generation model, PRISM-DP eliminates the need for manual mesh creation, improving scalability in open-set environments. Experiments in simulation and the real world show that PRISM-DP outperforms high-dimensional image-based policies and achieves performance comparable to policies trained with ground-truth state information.
I Introduction
@@
To address these challenges, we introduce Pose-tRacking via Image Segmentation and Mesh-generation Diffusion Policies (PRISM-DP), a framework that leverages segmentation, mesh generation, pose estimation, and pose tracking models to enable compact diffusion policy learning directly from the spatial poses of task-relevant objects. More specifically, PRISM-DP applies a segmentation model to isolate task-relevant objects in input images, it uses the resulting segmentation masks to automatically create 3D triangulated meshes for all target objects using a mesh generation model, then it uses the newly created meshes to initialize a pose estimation and tracking model. Subsequently, the diffusion policy is conditioned on both the robot state and the continuously estimated object poses for the target objects. Crucially, because PRISM-DP uses a mesh generation model, it eliminates the need for manual mesh processing or creation, improving scalability and usability in open-set, real-world environments.
@@
Through several experiments in both simulated and real-world settings, we demonstrate that PRISM-DP facilitates the learning of compact diffusion policies that achieve performance comparable to policies trained using ground-truth object pose information or manually generated meshes, while significantly surpassing image-based diffusion policies with similar parameter counts. To our knowledge, our work is the first to show that diffusion policies conditioned on pose-based observations can consistently outperform those conditioned on raw image observations in a real-world setting, suggesting exciting new directions for future research.
II Related Works
II-A 6D Pose Estimation and Tracking
@@ II Related Works
6D pose estimation and tracking are foundational technologies in robotics, with decades of extensive study and advancement [8]. Pose estimation refers to the process of determining the 3D position and orientation (collectively known as the 6D pose) of an object from sensor data, such as images or point clouds. Traditionally, pose estimation and tracking have been treated as distinct problems. Pose estimation methods are broadly categorized into model-based approaches, which rely on known CAD or mesh models of the object [16, 30], and model-free approaches, which instead use a few reference images of the target object [4]. Although model-free methods appear more user-friendly, they often require fine-tuning for each novel object [21], suffer from inefficiencies due to internal reconstruction processes [9], or fail under conditions of occlusion and low texture [10]. As a result, model-based methods are generally preferred for achieving robust and reliable pose estimation.
@@
Pose tracking, by contrast, focuses on continuously updating an object’s 6D pose over time by exploiting temporal information from sequential video frames. It is the task of estimating an object’s changing position and orientation across a sequence of frames, typically with an emphasis on efficiency and smoothness. Like pose estimation, tracking methods can be divided into model-free [38] and model-based approaches [32].
@@
II-B Policy Learning with Poses
Pose information provides a compact and structured description of the environment, making it ideal for policy learning. However, deploying pose-based policies in the real world remains challenging due to the difficulty of acquiring accurate 6D poses.
External tracking methods such as AprilTags [2] or Motion Capture [20] are commonly used but require specialized hardware and controlled environments. Others reconstruct scenes into point clouds or voxel maps using depth sensors [19, 40] or LiDAR [33, 34], but such methods are sensitive to environmental changes and require repeated preprocessing if changes occur. Diffusion policies initially trained with ground-truth poses in simulation have been adapted for real-world tasks via visual encoders that process input images [6, 36], but this technique comes at the cost of three to four times increased model size.
Recent work has incorporated FoundationPose [39] into learning-based robotics pipelines [14]. However, Pan et al. [25] use poses only for task planning, while Hsu et al. [14] requires manual mesh scans using consumer devices. In contrast, our approach removes the need for manual reconstruction by integrating segmentation and mesh generation for open-set 6D pose tracking.
II-C 3D Generation
The success of large-scale generative models for text and images has recently extended to 3D AI-Generated Content (3D AIGC), with diffusion [13] and Transformer-based architectures [37] driving progress in mesh generation. Recent methods fall into two categories: reconstruction-based approaches [3, 41], which first generate intermediate 3D representations before converting them into meshes, and autoregressive approaches [5, 31], which directly model mesh structures through sequence learning.
Reconstruction-based methods tend to produce higher visual fidelity [3, 41], while autoregressive methods generate meshes with more concise and elegant topology [5, 31] but often at the cost of visual quality [7]. Given the importance of geometric realism for downstream pose estimation and tracking, our approach uses a reconstruction-based mesh generation approach.
III Methodology
@@
(2) A segmentation model produces binary masks for each target object at their respective frames of first visibility:
$\mathbf{M}_{i}=\mathrm{Segmentation}(\mathbf{I}_{F_{i}}),\ \forall i\in\{1,...,J\}.$
Each pixel in a mask $\mathbf{M}_{i}\in\mathbb{R}^{H\times W}$ is $1$ if part of the $i$-th object at time $t$ and 0 otherwise.
(3) Each mask $\mathbf{M}_{i}$ is passed to a mesh generation model to obtain a triangular mesh:
$\mathcal{M}_{i}=\mathrm{MeshGen}(\mathbf{I}_{F_{i}}*\mathbf{M}_{i}),\forall i\in\{1,...,J\}.$
@@
The operation $\mathbf{I}_{F_{i}}*\mathbf{M}_{i}$ denotes elementwise multiplication, preserving RGB values in $\mathbf{I}_{F_{i}}$ where the mask is one and setting all other pixels to black (as seen in Fig. 1).
@@
(4) An approximate pose is calculated at the frames of first visibility for all target objects:
$\hat{\mathbf{T}}_{i}^{F_{i}}=\mathrm{PoseEstimator}(\mathbf{I}_{F_{i}},\mathbf{D}_{F_{i}},\mathcal{M}_{i},\mathbf{K}),\ \forall i\in\{1,...,J\}.$
(5) The pose estimates at the the frames of first visibility are used to bootstrap a pose tracking model per object:
$\begin{split}\hat{\mathbf{T}}_{i}^{t}&=\mathrm{PoseTracking}(\hat{\mathbf{T}}_{i}^{t-1},\mathbf{I}_{t},\mathbf{D}_{t},\mathcal{M}_{i},\mathbf{K}),\\ &\quad\forall i\in\{1,...,J\},\ \forall t\in\{F_{i},...,T\}.\end{split}$
@@
If an object is not visible at the beginning of a demonstration (i.e., $F_{i}\neq 1$ for some $i$), then the object poses at frames preceding $F_{i}$ are considered null, set as $4\times 4$ matrices of all zeros.
@@
The output of step 5 is a set of pose estimates over all task-relevant objects and time points. The pose estimates at time $t$ are concatenated with their corresponding robot state, $\mathbf{s}_{t}$, to create a sequence of observations, which are finally used to train a diffusion policy in a compact, low-dimensional space.
Conditioned on the observation history $\mathbf{o}_{t-H_{o}:t}$, the diffusion policy $\pi_{\theta}$ outputs a predicted sequence of actions $\mathbf{a}_{t:t+H_{p}}$. In a general diffusion model, a forward noising process gradually corrupts clean data $\mathbf{x}_{0}$ with Gaussian noise over a series of timesteps. Due to the Markov property of this process, this can be reparameterized into a closed-form solution that directly samples a noisy version $\mathbf{x}_{k}$ at any timestep $k$:
where $\boldsymbol{\varepsilon}$ is a standard Gaussian noise vector, and $\bar{\alpha}_{k}$ is a parameter from a predefined noise schedule that controls how much of the original signal $\mathbf{x}_{0}$ is preserved versus how much noise is added.
@@
Figure 4: Simulation tasks in Robosuite. Colored 3D bounding boxes indicate objects whose poses are estimated and tracked.
TABLE I: Success rates (SR) and training time per epoch (TE) across five Robosuite tasks. TE is recorded in seconds. Evaluated methods include Transformer-based (DP-T) and UNet-based (DP-U) diffusion policies trained with ground-truth object poses (GTP), raw RGB images (Img), and estimated 6D poses using ground-truth meshes (GTM). We also include larger image-based variants, DP-T-L Img and DP-U-L Img, that are 8 times and 11 times larger than their standard counterparts. Our proposed method with a Transformer and a UNet is denoted as PRISM-DP-T and PRISM-DP-U.
@@
During inference, the goal is to reverse this process. The model learns to gradually denoise a sample, starting from pure noise, through an iterative procedure. Specifically, at each timestep $k$, a learned denoising network $\epsilon_{\theta}(\mathbf{x}_{k},k,\mathbf{y})$ predicts the noise component $\boldsymbol{\varepsilon}$ present in the current noisy sample $\mathbf{x}_{k}$, where $\mathbf{y}$ represents conditioning information. Given this predicted noise, the model constructs the mean $\mu_{\theta}(\mathbf{x}_{k},k,\mathbf{y})$ of the reverse distribution used to sample the less-noisy version $\mathbf{x}_{k-1}$:
where $\alpha_{k}$ is another parameter derived from the noise schedule. The network $\epsilon_{\theta}$ is trained to predict the ground-truth noise $\boldsymbol{\varepsilon}$ by minimizing the mean squared error loss, where the expectation is taken over the data distribution, noise, and timestep:
$\mathcal{L}(\theta)=\mathbb{E}\left\|\boldsymbol{\varepsilon}-\epsilon_{\theta}(\mathbf{x}_{k},k,\mathbf{y})\right\|^{2}$
In PRISM-DP, we instantiate this diffusion model as our policy $\pi_{\theta}$ on action sequences. Therefore, the training objective therefore becomes:
$\mathcal{L}_{\text{PRISM}}(\theta)=\mathbb{E}\left\|\boldsymbol{\varepsilon}-\epsilon_{\theta}(\mathbf{a}_{t:t+H_{p},k},k,\mathbf{o}_{t-H_{o}:t})\right\|^{2},$
where we replace the generic data $\mathbf{x}_{k}$ with the noisy action sequence $\mathbf{a}_{t:t+H_{p},k}$ and the conditioning $\mathbf{y}$ with our observation window $\mathbf{o}_{t-H_{o}:t}$.
IV Evaluation
IV-A System Implementation
In this subsection, we provide implementation details for our system implementation that instantiates the approach outlined in §III. For the $\mathrm{Segmentation}$ subroutine, the base model in our system is Segment Anything 2 (SAM2) [28]. Task-relevant objects are specified to SAM2 by clicking on the target objects within the sequence of RGB images on their first visible frames, $\mathbf{I}_{F_{i}}$. Such inputs are called “point prompts”.
During task execution, our system can run at 10 Hz control frequency constantly, introducing no practical latency bottleneck. Initially, the poses of all task-relevant objects are set to null values (all zeros). When a target object first appears in the field of view, the user provides point prompts to SAM2. FoundationPose then uses the resulting segmentation mask and corresponding mesh object (cached during training) to initialize the object’s pose via pose estimation. To demonstrate the generalizability of our proposed system, all experiments are initialized from random states instead of in-distribution states.
IV-B Simulation Experiments
IV-B1 Experimental Settings
To evaluate the effectiveness of PRISM-DP, we conduct extensive experiments in simulation across five tasks from Robosuite [42], using a 7-DOF simulated Franka Panda arm. The control is executed in end-effector space, with rotations represented as quaternions. As illustrated in Fig. 4, the tasks are lifting a red cube (Lift), sorting a can into its designated compartment (Can), inserting a square nut onto a square peg (Square), stacking a red cube onto a green cube (Stack), and placing a mug into a drawer for cleanup (Mug).
The simulation evaluation includes a comprehensive set of baselines and architectural variants, as summarized in Table I. Evaluated methods include Transformer-based (DP-T) and UNet-based (DP-U) diffusion policies trained with ground-truth object poses (GTP), raw RGB images (Img), and estimated 6D poses using ground-truth meshes (GTM). We also include larger image-based variants, DP-T-L Img and DP-U-L Img, that are 8 times and 11 times larger than their standard counterparts (detailed further in the following paragraph). Our proposed method with a Transformer and a UNet is denoted as PRISM-DP-T and PRISM-DP-U.
IV-B2 Results
As shown in Table I, PRISM-DP consistently outperforms image-conditioned diffusion policies with the same network capacity while having better training efficiency, regardless of whether a Transformer or UNet architecture is used. It achieves comparable performance to diffusion policies trained with poses obtained from ground-truth meshes, indicating that learning-based mesh generation is sufficiently accurate to support downstream pose tracking and policy learning for these tasks. While scaling up the parameter count by 8 times and 11 times in the Transformer-based and UNet-based image models (DP-T-L Img and DP-U-L Img) closes the performance gap for simpler tasks such as Lift and Can, this improvement comes at a significant computational cost and fails to generalize to more challenging tasks. For instance, in tasks like Stack, Square, and Mug, PRISM-DP still outperforms these large image-based baselines, underscoring the value of low-dimensional structured observations.
@@ IV-B2 Results
Figure 5: Top left: Examples of ground-truth and generated meshes, shown in the upper and lower rows, respectively. Top right: Visualization of estimated and ground-truth poses of one object during a rollout. Bottom: Average PSNR, FID, and the average positional and orientation errors in radians across all objects in simulation training datasets.
To further understand the efficacy of PRISM-DP, we evaluate the quality of mesh generation as well as the accuracy of pose estimation and tracking. For mesh generation, we compare meshes produced by the generative model against ground-truth meshes from Robosuite. For pose estimation, we compare the predicted poses, estimated using the generated meshes, to the ground-truth poses transformed into the camera frame.
To assess mesh quality, we render both ground-truth and generated meshes from 30 randomly sampled viewpoints using Blender [1], and compute two commonly used perceptual metrics: Peak Signal-to-Noise Ratio (PSNR) [15] and Fréchet Inception Distance (FID) [12]. PSNR measures pixel-level reconstruction fidelity, where higher values indicate closer alignment, with 20–40 dB considered high quality [29]. FID assesses similarity in feature space, where lower scores are better, and values below 10 typically reflect high-quality generation [18].
@@
IV-C Real-World Experiments
IV-C1 Experimental Settings
To further validate the proposed approach, we conduct real-world experiments using a dual-arm robotic system (Fig. 6). The setup comprises two 7-DOF xArm7 manipulators, each mounted on a 1-DOF linear actuator. One arm is equipped with a RealSense D435i RGB–D camera for viewpoint control, while the other is fitted with a parallel gripper for manipulation. This configuration enables the policy to dynamically select viewpoints during task execution.